OpenStudy (anonymous):

Which of the following is a solution of x2 + 5x = -2?

OpenStudy (anonymous):

i can help :D

OpenStudy (anonymous):

it has to be a quadratic formula

OpenStudy (anonymous):

solution i meant

OpenStudy (anonymous):

no multiple choice?

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

i have it

OpenStudy (anonymous):

it'll be an image

OpenStudy (anonymous):

k

OpenStudy (mrnood):

first re-write the equation in 'standard' quadratic form: \[ax ^{2}+bx+c=0\]

OpenStudy (anonymous):

a=1 b=5 c=2

OpenStudy (anonymous):

x^2-5x+2=0

OpenStudy (mrnood):

ok - that's correct Now this does not factorise. SO you need to solve by another method Use the quadratic formula to get the two solutions to this equation: \[x=\frac{ -b \pm \sqrt{b ^{2}-4ac} }{ 2a }\]

OpenStudy (anonymous):

im trying to do the formula

OpenStudy (mrnood):

BTW - you mis typed the equation x^2+5x+2=0

OpenStudy (anonymous):

i got it -5+-sqroot 27 over 2

OpenStudy (anonymous):

correct?

OpenStudy (mrnood):

Yes correct BUT just be careful: the WHOLE thing is /2 (-5+- sqr(27))/2

OpenStudy (anonymous):

oh ok, next problem

OpenStudy (mrnood):

hold on b^-4ac = 5^2 -4(1*2) = 25-8 = 17

OpenStudy (anonymous):

wasn't that i put down

OpenStudy (anonymous):

so may i do next problem

OpenStudy (anonymous):

Solve 3x2 + 4x = 2.

OpenStudy (anonymous):

i don't understand the choices they have

OpenStudy (anonymous):

um, nvm, i sorta had a blank mind and the answer was right there

OpenStudy (mathmale):

1) Please post each new problem separately. 2) More important than understanding the answer choices is that you yourself learn how to solve these problems. 3) The first step towards solving 3x2 + 4x = 2 is to move that 2 to the left side of this equation. To accomplish that, subtract 2 from both sides. doing this puts your equation into the required standard form: y=ax^2+bx+c=0

OpenStudy (anonymous):

oh ok thx

OpenStudy (mathmale):

then you have several choices of approach: quadratic formula, factoring, graphing.

OpenStudy (anonymous):

i did the quadratic formula, and the answer i got was -2+- 2sqrt 10 over 3

OpenStudy (anonymous):

is that correct?

OpenStudy (mathmale):

Now please check your answer by substituting it back into the original equation.

OpenStudy (anonymous):

how do i do that

OpenStudy (mathmale):

I'm sure you've done this kind of check before. Supposing that you started with the equation 3x2 + 4x = 2 and that you came up with the possible answer x=1. You'd substitute x=1 into this equation: 3(1)^2+4(1)=2. Is this equation true or false?

OpenStudy (anonymous):

false

OpenStudy (anonymous):

7 doesn't equal 2

OpenStudy (mathmale):

Right. That's not important; what is important is that you be able to check your possible solutions by substituting them back into the original equation.

OpenStudy (mathmale):

'i did the quadratic formula, and the answer i got was -2+- 2sqrt 10 over 3'

OpenStudy (mathmale):

What's your next step?

OpenStudy (anonymous):

idk...

OpenStudy (mathmale):

What did I just ask you to do, as an example?

OpenStudy (anonymous):

just substitute

OpenStudy (anonymous):

and i got it wrong

OpenStudy (mathmale):

The whole point, Minias, is that if you want to check your possible roots, you'll need to substitute them back into the original equation. Have you tried that with your possible roots \[x=-2+- 2\sqrt 10 \over 3\] or \[x=\frac{ -2\pm \sqrt{10} }{ 3 }?\]

OpenStudy (mathmale):

Your 'x2 + 5x = -2' becomes \[1x^2+5x+2=0\]

OpenStudy (anonymous):

oh

OpenStudy (mathmale):

...and from this you can see that the coefficients 'a', 'b' and 'c' are 1, 5 and 2.

OpenStudy (anonymous):

ahh i think i somewhat got it

OpenStudy (mathmale):

I'm sorry, but I was focusing on the wrong problem. This is why I asked you to post each new problem separately. Am I correct in believing that we're now focused on solving 3x2 + 4x = 2?

OpenStudy (mathmale):

Please subtract 2 from both sides of this equation (assuming that you want to continue).

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

3x2+4x-2

OpenStudy (mathmale):

You've started with an equation and must end with an equation. Please type "=0" after your 3x^2+4x-2.

OpenStudy (anonymous):

3x^2+4x-2=0

OpenStudy (mathmale):

Good. Now, identify the coefficients a, b and c.

OpenStudy (anonymous):

a=3 b=4 c=-2

OpenStudy (mathmale):

Good. Now take those three coeff. values and substitute them into \[x=\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\]

OpenStudy (anonymous):

how do i make that?

OpenStudy (mathmale):

All I'm asking you to do is to substitute a=3, b=4 and c=-2 into the quadratic formula.

OpenStudy (mathmale):

Hint: b=4, so we get\[x=\frac{ -4\pm \sqrt{4^2-4(a)(c)} }{ 2a }\]

OpenStudy (mathmale):

substitute a=3 and b=-2 into this formula, please.

OpenStudy (anonymous):

-4(3)(-2)

OpenStudy (anonymous):

which would be +24

OpenStudy (mathmale):

Yes, that's part of the discriminant, \[b^2-4ac,\], but you'll need to complete all of the required substitutions into \[x=\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\]

OpenStudy (mathmale):

Again, if b=4, we get \[x=\frac{ -4\pm \sqrt{4^2-4(a)(c)} }{ 2a }\]

OpenStudy (mathmale):

Please substitute a=3 and c=-2 into this formula.

OpenStudy (anonymous):

4(3)(-2) 2(3)

OpenStudy (mathmale):

Note: 4(3)(-2)=12(-2) = -24. Where did your 2(3) come from?

OpenStudy (anonymous):

3 =a

OpenStudy (mathmale):

I'm asking you to start with \[x=\frac{ -4\pm \sqrt{4^2-4(a)(c)} }{ 2a }\] and to substitute the values a=3 and b=-2. Your answer must look like\[x=\frac{ -4\pm \sqrt{4^2-4(a)(c)} }{ 2a }\] or it'd not be complete.

OpenStudy (anonymous):

sry i meant -4

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

\[\frac{-4\pm \sqrt{4^2-4(3)(-2)} }{ 2(3) }\]

OpenStudy (anonymous):

like that?

OpenStudy (mathmale):

Nice! and that reduces to\[x=\frac{-4\pm \sqrt{16+24} }{ 6 }\]

OpenStudy (mathmale):

can you simplify this further?

OpenStudy (anonymous):

\[\frac{ -4\pm \sqrt{40} }{ 6}\]

OpenStudy (mathmale):

Very good. Be sure to label this:\[x=\frac{ -4\pm \sqrt{40} }{ 6}\]

OpenStudy (mathmale):

and to reduce it. Note that Sqrt(40)=Sqrt(4)*Sqrt(10)=2Sqrt(10).

OpenStudy (anonymous):

wt would 40 reduce to?

OpenStudy (mathmale):

reduce it now, please:\[x=\frac{ -4\pm \sqrt{40} }{ 6}=\frac{ -4\pm2\sqrt{10} }{ 6 }=?\]

OpenStudy (anonymous):

nvm

OpenStudy (mathmale):

40 factors into 4 and 10, and 4 is a perfect square, so Sqrt(40)=2Sqrt(10).

OpenStudy (anonymous):

\[\frac{ 4\pm \sqrt{10} }{ 3 }\]

OpenStudy (anonymous):

-4 i meant

OpenStudy (mathmale):

Please take another look at that:\[\frac{ -4\pm2\sqrt{10} }{ 6 }=\frac{ -4 }{ 6 }\pm \frac{ 2\sqrt{10} }{ 6 }=?\]

OpenStudy (mathmale):

-4/6 reduces to -2/3. 2Sqrt(10)/6 reduces to what?

OpenStudy (anonymous):

\[\frac{ -2\pm \sqrt{10} }{ 3}\]

OpenStudy (mathmale):

Good. That's it. Any questions? Be sure to label your results.\[x=\frac{ -2\pm \sqrt{10} }{ 3}\]

OpenStudy (anonymous):

one more

OpenStudy (anonymous):

but i'll make another post on it

OpenStudy (mathmale):

Please take a moment to think about what you've learned from solving these two problems. I'd suggest you take notes so that you have something to which to refer next time you encounter a problem requiring you to find the roots of a quadratic equation.

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

I got all of them right, thx

OpenStudy (mathmale):

Happy to hear that.